A light source emits a photon during a time interval of 1.87 10-8 s.

(a) Find the minimum uncertainty in the photon's energy.

(b) Find the minimum frequency bandwidth of the light.
Answer #1

a) delta E*deltat = h/4pi

delta E = 6.63*10^-34/4*3.14*1.87*10^-8 = 0.28*10^-26 J

b) delta E = hf

f = 0.28*10^-26/6.63*10^-34 = 0.0422*10^8 Hz